Solve the linear and inequalities
1)3s-1÷4+s+2÷3=3÷2
2)3(x-2)-2<3-(x+4)÷1

Answer by Amily_2190(27)

Use PEMDAS
1)
3s-1÷4+s+2÷3=3÷2
3s-(1/4)+s+(2/3)=(3/2)
4s+(5/12)=(3/2)
(subtract 5/12 from both sides to isolate 4s)
4s=(13/12)
s= (13/12)/4= (13/48)

s=(13/48)

2)3(x-2)-2<3-(x+4)÷1
3x-6-2<3-(x+4)
3x-8<3-x-4
3x-8<-1-x
(Add x to both sides because of inverse operation)
4x-8<-1
(Add 8 to both sides; inverse operation)
4x<7
(Divide both sides by 4; inverse operation)
x<7/4